OpenStudy (anonymous):

One leg of the right triangle is 20cm. and the hypotenuse is 10 cm. longer than the other leg. Find the lengths of the unknown sides?

OpenStudy (anonymous):

\[10\sqrt{?}\]

OpenStudy (anonymous):

10sqrt5

OpenStudy (anonymous):

i need solution please

OpenStudy (anonymous):

solution please..

OpenStudy (anonymous):

(20+10)^2=20^2+b^2 30^2-20^2=b^2 900-400=b^2 10sqrt5=b

OpenStudy (anonymous):

thanks..

OpenStudy (anonymous):

c^2=a^2+b^2 c=hypotenuse,a=other side,b=20 (a+10)^2= a^2+(20)^2 a^2+20a+100=a^2+400 a^2-a^2+20a=400-100 20a=300 a=300/20=15 so length of unknown side is 15 and hypotenuse=15+10=25

OpenStudy (anonymous):

soxh is wrong

OpenStudy (anonymous):

OMG thanks

OpenStudy (anonymous):

25 is the final answer

OpenStudy (anonymous):

hypotenuse=25 and other side=15

OpenStudy (anonymous):

can i aski another question

OpenStudy (anonymous):

A and B start from a given point and travel on a straight road at average speed of 40 and 60 kph respectively. If be starts 3 hours after A, find the time and the distance they travel before meeting.

OpenStudy (anonymous):

A and B start from a given point and travel on a straight road at average speed of 40 and 60 kph respectively. If be starts 3 hours after A, find the time and the distance they travel before meeting.

OpenStudy (anonymous):

time_to_meet_after_b_starts=(3h*40km/h)/(60km/h-40km/h)=120km/(20km/h)=6h time_to_meet_after_a_starts=3h+6h=9h distance = 9h*40km/h = 360km, can alternatively be calculated also dsitance = 6h * 60km/h = 360km